# RAGBench: Explainable Benchmark for Retrieval-Augmented Generation Systems

### v1

@@ 1 Introduction
We evaluate state-of-the art LLMs and existing RAG evaluation systems on RAGBench. We find that, while few-shot LLM judges perform equally well across domains and task types, they still under-perform compared to a fine-tuned DeBERTa-large model. We motivate future work to leverage these data for advancing RAG evaluation approaches and improve on the proposed benchmark.
@@ 2 Related Work
RAG benchmarks
Numerous general LLM evaluation benchmarks, such as ChatbotArena [46] have been proposed in past work. However, human preference datasets, constructed through pairwise comparisons, have limitations. While these data are appropriate for fine-tuning general purpose LLM judges, they are insufficient for building RAG evaluation systems because preference judgements under-represent important RAG dimensions like factuality and completeness of the response [13].
ChatRAGBench [24] is a recent initiative that is similar in intent to our work in that it contributes a large-scale unified RAG benchmark. However, ChatRAGBench only contains ground truth responses and lacks the granular component-specific labels that we release with RAGBench. As future work, we can consider annotating ChatRAGBench with the schema proposed in this paper, to further scale RAGBench.
RAGTruth [41] is another recent effort at a RAG Benchmark. RAGTruth combines QA, Data-toText, and Summarization RAG data with human annotated hallucinated spans in the response. While it is an excellent benchmark for hallucination detection, it does not offer the level of granularity we present with RAGBench that is necessary to understand the RAG system as a whole.
@@
Recently, several parallel efforts have proposed approaches to automated RAG evaluation. In RAGAS [9], the authors query an LLM-judge (GPT-3.5) with a curated prompt to evaluate context relevance, answer relevance and faithfulness of a RAG response. Next, Saad-Falcon et al. [33] propose ARES, a framework for fine-tuning smaller NLI models to predict the same metrics. This approach benefits from fine-tuning, though domain-specific annotated validation sets are required for each domain adaptation. In parallel, Chen et al. [4] develop a heuristic system to probe LLM’s robustness to noisy and irrelevant context documents, and Adlakha et al. [1] explore heuristic algorithms to estimate RAG correctness and faithfulness. The lack of established RAG benchmarks makes it difficult to compare these approaches against each other. We aim to address this limitation by introducing RAGBench.
@@ 3.3 LLM annotator
For high quality labels, we use proven techniques like chain of thought [40] that have been shown to maximize the correlation between GPT-4 and human judgements [43, 46]. For relevance and utilization we request the LLM-annotator to directly identify relevant and utilized sub-strings in the input documents. For adherence, we instruct the LLM to identify which response sentences, if any, are supported by the provided context. We can then derive an example-level boolean adherence label by checking if all response sentences are supported. The exact prompt used for annotation is provided in Appendix 9.4. We apply post-processing steps to ensure high quality, reliable annotations from our GPT-labeler, which we outline in Appendix 9.5. We further validate our annotation approach in Section 4, and discuss the limitations of using an LLM-annotator in Section 8.
RAGBench raw annotations contain token-level labels for utilization and relevance, which are converted to TRACe metrics using equations in Section 3.2. We encourage future work on automated evaluators to predict the raw token-level labels, like relevant and utilized spans, rather than predicting the example-level scores directly which are less interpretable for the end user.
Table 2: Ranking of Simulated RAG Systems. We evaluate GPT-4-turbo annotations on simulated RAG datasets from Saad-Falcon et al. [33]. The data from each source are synthetically augmented to create sets with increasing degrees of context relevance (Rel) and answer adherence (Adh). We annotate 500 samples from each set and rank them according to the average context relevance and answer adherence metrics. We report Kendall’s tau to evaluate the agreement between GPT-4-turbo rankings and ground truth (higher is better).
NQ HotpotQA WoW FEVER
Rel Adh Rel Adh Rel Adh Rel Adh
Kendall’s Tau binary 1.0 0.83 0.87 1.0 1.0 0.89 1.0 0.78
Kendall’s Tau continuous 0.94 - 0.73 - 1.0 - 0.77 -
4 Annotation Validation
We validate out metric formulations and labeling approach on simulated RAG datasets of varying quality. We use mock RAG datasets generated by Saad-Falcon et al. [33] for this analysis. Their RAG validation set is sampled from KILT [30], including Natural Questions (NQ)[18], HotpotQA[42], FEVER[36], and Wizards of Wikipedia (WoW) [8] datasets. The authors synthetically generate systems of varying quality by adjusting the ratio of relevant documents and responses in the data. We sample 500 examples from each simulated RAG dataset and annotated them as described in section 3.3. Next, we calculate average annotated context relevance and adherence scores for each dataset and use those to rank the mock systems. We compare our rankings to ground truth with the Kendall rank correlation (Kendall’s $\tau$) metric, which evaluates the agreement between two sets of ranks on a scale from 0 (no agreement) to 1 (perfect agreement).
As shown in Table 2, the GPT-4 annotations achieve high Kendall’s $\tau$ ranging from 0.78 to 1. For a fair comparison with the ground truth labels, we derive binary context relevance and labels from the GPT-4 annotations by thresholding the example Relevance score (equation 2) at 0. For comparison, we also report ranking results with out more granular example-level Relevance scores that range from 0-1. We find that these metric produce a different ranking (see lower Kendall’s $\tau$ in Table 2), which we attribute to the metrics capturing differences in retrieved context length across the different examples.
@@ 6 Discussion
LLMs underperform on the RAG evaluation task
We observe that the finetuned DeBERTa model outperforms the few/zero-shot LLM-judge baselines on most datasets. While GPT-3.5 demonstrates competitive performance with DeBERTa on a few metrics, DeBERTa consistently achieves superior performance metrics across all evaluations. Despite the versatility of LLM judges across various tasks, their lack of specialization necessitates finetuning for optimal results. Future work may focus on finetuning LLM judges to close the gap between DeBERTA and GPT-4 evaluation performance. In Appendix 9.7, we demonstrate that, despite its small size, the finetuned DeBERTA model does generalize to out of domain RAG datasets in the same way that LLM-based approaches do.
@@ 7 Conclusion
In this paper we introduce RAGBench, a large-scale dataset composed of real-world RAG examples intended for training and benchmarking RAG evaluation models. Additionally, we formulate TRACe, a RAG evaluation framework comprising four metrics: uTilization, Relevance, Adherence, and Completeness. TRACe standardizes the evaluation process, offering a consistent and systematic approach to measuring RAG system performance across various dimensions.
We benchmark existing RAG evaluation framework using RAGBench and demonstrate that LLM-judges struggle to compete with a fine-tuned RAG evaluation expert model. Future work may involve fine-tuning larger expert models to explore the potential for narrowing the performance gap between these models and the ground truth.
Our contributions address the need for standardized benchmarks and methodologies, enabling more precise and actionable insights into the strengths and weaknesses of different RAG systems. This, in turn, will facilitate the iterative improvement of RAG models, driving forward the capabilities of retrieval-augmented generation in real-world applications.
8 Limitations
LLM Annotations
Though LLMs demonstrate high correlations with human judgements on a variety of tasks [6, 11], using them as a singular source of ground truth remains controversial [23]. At the same time, human judgements of LLM outputs are also prone to inconsistencies and bias. In [13], the authors find that human evaluators are often misled by the assertiveness and complexity of the LLM model output, which leads them to underestimate the rate of factuality errors in LLM responses.
In this work, we acknowledge the potential of noise and bias in RAGBench resulting from automated GPT-4-turbo annotations, and the concerns about the potential transmission of such biases into subsequent RAG systems. One way to address this in future may be to replace the GPT-4-annotator with and LLM "jury" as suggested in [38]. By aggregating judgements from diverse models, this approach can help reduce the noise and bias in the output judgements at low cost.

### v2

@@ 1 Introduction
We evaluate state-of-the art LLMs and existing RAG evaluation systems on RAGBench. We find that a 400M-parameter DeBERTa-large model that was fine-tuned on RAGBench outperforms few-shot LLM judges across numerous domains and task types. We highlight this result to motivate future work aimed at leveraging these data for advancing RAG evaluation models and improving on the proposed benchmark.
@@ 2 Related Work
We differentiate our work from existing ground-truth RAG datasets like ChatRAGBench [23], CRAG [41], DomainRAG [38], and ALCE [10]. These datasets contain RAG samples with ground-truth responses and are used for end-to-end evaluation of RAG systems via response-level metrics like exact match or ROUGE scores. In contrast, we design RAGBench to enable development of more mature evaluation systems that effectively evaluate different parts of the RAG system along multiple dimensions like retriever relevance, adherence and completeness of the response.
Various RAG benchmarks focus specifically on hallucination detection [40, 33, 21, 5]. FELM [5] is a multi-domain and task dataset with factuality labels for open domain QA. RAGTruth [40], DelucionQA [33], and HaluEval [21] are RAG-specific datasets with both synthetic as well as human-annotated labels for hallucinations in LLM reseponses. While these are appropriate benchmarks for hallucination detection models, they do not offer the level of granularity we offer with RAGBench that is necessary to understand the RAG system as a whole.
@@
Recently, several parallel efforts have proposed approaches to automated RAG evaluation. In RAGAS [9], the authors query an LLM-judge (GPT-3.5) with a curated prompt to evaluate context relevance, answer relevance and faithfulness of a RAG response. Next, Saad-Falcon et al. [32] propose ARES, a framework for fine-tuning smaller NLI models to predict the same metrics. In parallel, Chen et al. [4] develop a heuristic system to probe LLM’s robustness to noisy and irrelevant context documents, and Adlakha et al. [1] explore heuristic algorithms to estimate RAG correctness and faithfulness. The lack of established RAG benchmarks makes it difficult to compare these approaches against each other. We aim to address this limitation by introducing RAGBench.
@@ 3.4 LLM annotator
For high quality labels, we use proven techniques like chain of thought [39] that have been shown to maximize the correlation between GPT-4 and human judgements [43, 46]. For relevance and utilization we request the LLM-annotator to directly identify relevant and utilized sub-strings in the input documents. For adherence, we instruct the LLM to identify which response sentences, if any, are supported by the provided context. We can then derive an example-level boolean adherence label by checking if all response sentences are supported. The exact prompt used for annotation is provided in Appendix 7.4. We apply post-processing steps to ensure high quality, reliable annotations from our GPT-labeler, which we outline in Appendix 7.5.
Table 2: GPT-4 annotator achieves high alignment with human judgements. We report F1 and Accuracy alignment metrics on human annotated subsets of DelucionQA. Adherence annotations are evaluated against DelucionQA human-annotated labels. Relevance and Utilization are evaluated against DelucionQA(40): a subset of 40 examples randomly sampled from the DelucionQA test set and annotated by the authors.
Test Set Metric F1 Accuracy
DelucionQA - example level Adherence 0.96 0.93
DelucionQA - span level Adherence 0.97 0.95
DelucionQA(40) - span level Utilization 0.92 0.94
DelucionQA(40) - span level Relevance 0.76 0.78
Alignment with Human Judgements
We validate our metric formulations and labeling approach on a human annotated benchmark. DelucionQA [33] is a curated collection of user queries on the operation of Jeep’s 2023 Gladiator model. Natural language queries are first generated by an LLM, then reviewed and filtered by human annotators. Context documents are sourced from Jeep’s Gladiator User Manual, and responses are generated by various LLMs. Combined with example-level and span-level annotations for hallucination, DelucionQA represents a realistic distribution of real-world user queries and RAG responses. We find that our GPT annotator achieves 93% and 95% example- and span-level agreement with human judgements on the DelucionQA test split (Table 2).
To validate relevance and utilization annotations, we also annotate a small subset of DelucionQA with granular relevance and utilization labels. We refer to this subset as DelucionQA(40) in Table 2. Similar to adherence, we observe high correlation between relevance and utilization judgements from GPT-4 and humans. Details of the annotation process and additional validations are found in Appendix 7.6.
(a) Retriever vs. Relevance
(b) LLM vs. Hallucination
(c) Prompt vs. Utilization
(d) Prompt vs. Completeness
Figure 4: Relationship between RAG system configuration and TRACe metrics. (a) The choice and configuration of RAG retriever component affects the average relevance of the retrieved context. In this example, a dense retriever with a low number of documents per query (k=2) yields the highest average context relevance. (b) The choice of LLM and generation prompt affect how well the RAG system utilizes the provided context. Prompting the LLM with a detailed chain-of-thought prompt leads to reduced hallucinations, as well as higher response utilization (c) and completeness (d) rates.
3.5 RAG Case Study
We design a case study to further motivate and validate the proposed TRACe framework. We sample 100 realistic world knowledge queries from the RAGTruth [40] QA training set, along with 2,512 unique context document chunks from the same dataset to use as inputs into our mock RAG systems. We simulate RAG setups of varying quality by controlling four configurable parameters: the retriever (sparse TF-IDF vs. dense), number of retrieved documents (2 vs. 4), generation model (GPT-3.5-turbo vs. GPT-4o), and 4 versions of the generation prompt template. For illustrative purposes, we evaluate one prompt template comprising of only the question (no context) vs. three RAG-style templates that encourage the LLM to utilize the provided context and/or utilize chain-of-thought (CoT). The full generation templates are provided in Appendix 7.7. We generate responses for the 100 RAGTruth queries with each of the 32 resulting RAG systems and use our LLM annotation prompt to evaluate the TRACe metrics.
Figure 4 demonstrates how the different RAG configurations influence TRACe metrics. For example, we confirm that the choice of the generative LLM model affects the amount of hallucinations (or non-adherent responses) in the generated text. Not surprisingly, GPT-4o combined with a chain-of-thought prompt yields the lowest amount of hallucination compared to the weaker GPT-3.5 model and less detailed prompts. Curiously, we find that GPT-3.5 hallucinates more when prompted with CoT, which may point to limitations in the model’s ability to reason about complex concepts 4(b). Overall, we find that prompting the LLM to think step-by-step and explain its reasoning leads to higher utilization of the provided context and more complete responses (4(c), 4(d)). Finally, we show that the choice of the retriever affects average relevance of the retrieved context documents per query 4(a).
@@ 5 Results
We observe that the finetuned DeBERTa model trained on RAGBench achieves competitive performance with billion-parameter LLM judges across numerous domain subsets of the RAGBench test set. On the hallucination detection task, DeBERTA AUROC scores range from 0.64 to 0.86. While RMSE for relevance and utilization range from 0.04 to 0.26, depending on the domain and task.
@@ 6 Conclusion
In this paper we introduce RAGBench, a large-scale dataset composed of real-world RAG examples intended for training and benchmarking powerful RAG evaluation models. To this end, we formulate TRACe, a RAG evaluation framework comprising four metrics: uTilization, Relevance, Adherence, and Completeness. TRACe standardizes the evaluation process, offering a consistent and systematic approach to measuring RAG system performance across various dimensions. We propose an automated approach to generate TRACe labels for RAGBench with an LLM and demonstrate high correlation between our approach and human judgements.
We benchmark existing RAG evaluation frameworks on RAGBench and demonstrate that a 400M parameter DeBERTa model finetuned on RAGBench data performs competitively with billion-parameter LLM Judges and commercial RAG evaluation systems. Though, the gap between the best-performing RAG evaluator and ground truth is still large. We motivate future work to leverage RAGBench toward fine-tuning more powerful evaluation models to explore the potential for narrowing the performance gap between these models and the ground truth.
Our contributions address the need for standardized benchmarks and methodologies, enabling more precise and actionable insights into the strengths and weaknesses of different RAG systems. This, in turn, will facilitate iterative improvement of RAG models, driving forward the capabilities of retrieval-augmented generation in real-world applications.